# Practical X-ray Gastric Cancer Screening Using Refined Stochastic Data Augmentation and Hard Boundary Box Training

### v4

@@ 5 Results
Figure 5 shows an example of the augmented image obtained with the proposed R-sGAIA for the leftmost original image. Figure 6 compares cancer detection on the EfficientDet-D7 network (baseline model 2) with and without the proposed R-sGAIA and HBBT. The input images are shown in (a), the detection results with RandAugment in (b), and those with the proposed R-sGAIA+HBBT in (c). The proposed system (c) accurately detected the gastric cancer region without being affected by mucosal irregularities or small areas, as demonstrated in the top and middle rows. However, as shown in the bottom row, the result nearly covers the ROI but is slightly misaligned, leading to a false-positive evaluation.
The detection performance of gastric cancer regions from gastric X-ray images is summarized in Table 2. The object detection model detects regions where the probability (i.e., the confidence score) of belonging to the class of interest exceeds a predefined detection threshold of the bounding box $\xi$ as locations containing objects of that class. HBBT assigns new negative labels to the over-detected regions that exceed the threshold $\xi$ and retrains the model. Therefore, to draw an ROC curve for evaluation, the model must be trained every time $\xi$ is changed, which is expensive. In this experiment, therefore, the detection threshold $\xi$ was determined such that the SE (= recall) of detecting gastric cancer was about 90%, which was higher than the doctor’s 85.5%. At this detection threshold, HBBT improves precision by 2.3 points in Faster R-CNN (baseline model 1) and 3.3 points in EfficientDet-D7 (baseline model 2). Our system, using the proposed techniques, R-sGAIA and HBBT, achieved a recall of 90.2%, a precision of 42.5%, and an F1-score of 0.578 (Table 2).
The combination of the proposed R-sGAIA and HBBT showed a significant improvement over RandAugment, one of the most advanced data augmentations, in the detection of gastric cancer using the general object detection model. Specifically, the improvement was 10.3 points and 6.1 points in precision, and 10.9 points and 5.9 points in F1 scores for Faster R-CNN (model 1) and EfficientDet-D7 (model 2), respectively.
The processing time required for one X-ray image is approximately 0.51 seconds, making the proposed system a practical and efficient screening tool for gastric X-ray images.
Figure 6: Example of cancer detection on EfficientDet-D7 networks (as shown in red box): (a) ground-truth region, (b) results obtained with RandAugment (baseline model 2), and (c) those with R-sGAIA+HBBT (proposed). The proposed R-sGAIA+HBBT led to the accurate detection for the top and middle examples, whereas for the bottom result, although it showed some improvement, the detection was inaccurate.
@@
Augmentation HBBT Precision Recall F1-score
RandAug sGAIA R-sGAIA
@@ 6 Discussion
The cancer detection performance of the proposed system in this study was significantly better than that achieved using baseline methods, including RandAugment (a state-of-the-art general data augmentation strategy), existing methods such as sGAIA Okamoto et al. (2019), and those proposed by (Laddha et al., 2019). Additionally, the system’s cancer detection capability exceeded that of a physician’s reading ($\sim$85.5%). In more than 2 out of 5 cases, the detection results presented by the system were cancerous. For a physician using the system, this means that by simply checking the areas surrounded by the few bounding boxes presented in a gastric X-ray image, a higher diagnostic yield can be achieved than with a conventional reading. In other words, the system is practically effective for on-site medical examinations and can significantly reduce the burden on physicians.
@@ 6.1 R-sGAIA
The proposed R-sGAIA can probabilistically highlight the gastric folds on a pixel-by-pixel basis based on medical findings, as shown in Figure 5. In this study, we found that R-sGAIA achieved an F1-score improvement of 6.0 points and 2.7 points over RandAugment on baseline models 1 and 2, respectively. This highlights the importance of probabilistic augmentation informed by medical knowledge. In baseline model 2, R-sGAIA further improved the F1-score by 1.2 points over the previous sGAIA in a direct comparison. This improvement may be attributed to R-sGAIA’s use of validation data to determine $p(e)$ as a continuous function, enabling it to generate a wider variety of enhancement images that better highlighted the gastric folds compared to sGAIA, which used a subjectively fixed value. The function of $p(e)$ with respect to edge strength $e=E(x,y)$ ranged approximately between 0.1 and 0.9, which likely contributed to the performance improvement. By allowing more stochastic variation in regions with very strong and weak edge strengths, R-sGAIA enhanced the diversity of the generated images, leading to better detection performance.
@@ 6.2 HBBT
Recent state-of-the-art object detection models (e.g., YOLOv7 (Wang et al., 2023)) are trained to minimize the difference between the boxes specified in the training data and the currently detected boxes, both in terms of their classes (i.e., contents) and locations. In other words, the model’s error is the sum of the class error and position error for each box, and training is performed to minimize these errors.
The proposed HBBT introduces a new and effective strategy that allows learning from healthy control images, which cannot be used to train conventional object detection models. It retrains over-detected boxes by assigning them a unique “hard-sample” class label to suppress future false detections. The HBBT strategy increases the class error for error-prone regions (i.e., hard samples) and retrains the model to reduce false positives. In essence, the HBBT learning algorithm is a strategy to reduce generic errors in object detection models, rather than focusing on the specific type of model. Furthermore, HBBT actively incorporates negative samples, such as images of healthy controls, which are typically excluded from training by conventional object detection models.
Although HBBT is a simple method that recursively learns to eliminate false positives detected in images, including control images, it alone improves precision by 2.3 points and 3.3 points and F1-score by 4.9 points and 2.9 points on baseline models 1 and 2, respectively. HBBT is a general-purpose learning method that can be combined with any object detection algorithm, not just for gastric cancer detection, as demonstrated in this study. Since the final model detection is determined by the bounding box confidence threshold, HBBT not only reduces the number of false positives but also helps minimize false negatives (as seen in the middle case in Figure 6).
6.3 For practical screening
In this diagnostic support system, the ability to handle the diversity in size and shape of the tumor area to be detected depends on the detectability of the object detection model used. The range of object sizes that can be detected is determined by hyperparameters such as anchor size and aspect ratio. In the task of cancer detection from gastric X-ray images, the general settings of the object recognition model are sufficient, as the images are taken in a way that ensures the tumor size is large enough for visual diagnosis by a physician. However, for extremely small tumors, it is necessary to adjust hyperparameters to account for their size.
@@ 6.3 For practical screening
Detecting gastric cancer regions from X-ray images is inherently challenging, not only for automatic diagnostic systems but also for experts. Despite these challenges, our two technological innovations, R-sGAIA and HBBT, have significantly improved the detection performance of gastric cancer, achieving a level of performance and speed that can be effectively used by physicians as a diagnostic aid. Our diagnostic support system offers a concrete solution to the challenges of gastric radiography, which has been easy to acquire images but difficult to diagnose. In the future, we aim to work closely with experts to further refine and improve the system.
@@ 6.4 Comparison with automated diagnostic support methods for other modalities
Many diagnostic support studies have focused on endoscopy, the most commonly used method for diagnosing gastric cancer. However, as mentioned earlier, most of these studies did not properly separate the training data from the evaluation data, leading to data leakage and artificially inflated diagnostic performance. Recent studies that have addressed this issue by using more sophisticated machine learning methods (Shibata et al., 2020; Ahmad et al., 2023) have reported detection performance for gastric cancer with an image-based F1 score of approximately 0.7, which is considered the current standard for diagnostic accuracy.
Due to its noninvasiveness and ability to image the entire body, X-ray computed tomography (CT) is widely used for detecting not only gastric cancer but also other organs and diseases. Diagnostic studies of gastric cancer using CT typically rely on classical radiomics, including necessary pre-treatment steps. Depending on the cancer stage, diagnostic performance usually achieves an AUC of 0.7 to 0.8 (Meng et al., 2020; Huang et al., 2022). The application of deep learning in this context primarily focuses on predicting patient prognosis and segmenting tumor areas rather than directly diagnosing cancer. Hao et al. (2022) demonstrated that radiomic features and machine learning-extracted features (i.e., low-dimensional representations) are effective in predicting the prognosis of gastric cancer patients, indicating that these techniques will become more prevalent in the future.
While the diagnostic performance of these automated support methods may appear lower than the 95% SE reported by physicians for gastric cancer, it is essential to note that this comparison is patient-based, making direct comparisons challenging. As automated diagnostic methods are introduced into clinical practice, their actual performance is likely to approach the current SE of physicians, particularly when multiple outputs are integrated to inform a final decision, providing significant support to physicians. Although our F1 score of 0.578 for the diagnostic performance of gastric X-ray images may seem lower than that of other modalities, it is important to recognize that this is a box-based evaluation. Additionally, since physicians diagnose gastric cancer by reviewing multiple images during a gastric X-ray examination, our method’s ability to highlight the ROI in each image makes it highly valuable. Furthermore, our approach is significant as it represents the first use of gastric X-ray images to aid in the diagnosis of gastric cancer.
@@ 6.5 Limitations of the study
In recent years, it has become apparent that medical images, such as MR images, contain features unique to each facility due to differences in imaging equipment, protocols, geometry, and other factors. These so-called domain differences can significantly impact the results of machine learning tasks, such as classification, when applied without adjustment. The gastric X-ray images used in this study were exclusively taken at medical facilities affiliated with Tokai University. While domain differences may affect the analysis of gastric X-ray images obtained from different facilities, this has not been verified in this study.
@@ 7 Conclusions
In this paper, we have proposed an unprecedented, accurate, and fast gastric cancer screening system from gastric X-ray images, built on two key innovations: (1) R-sGAIA and (2) HBBT. The proposed system (R-sGAIA + HBBT on the EfficientDet-D7 network) achieves a SE of 90.2%, surpassing that of physicians (85.5%), with a processing time of approximately 0.5 seconds per image. Notably, 42.5% of the detected candidate box regions contain cancer, meaning our system effectively narrows down the areas that physicians need to examine. This system also achieves an F1 score 5.9 points higher than the same network using RandAugment, a state-of-the-art data augmentation method. Diagnosing gastric cancer using gastric X-rays is a challenging task that requires extensive experience. We hope that our proposal will assist in diagnosis, leading to early detection and a reduction in the number of deaths in the near future.
@@ Acknowledgments
H. Okamoto was responsible for the practical system development of the study and the writing of the paper; T. Nomura, K. Nabeshima, and J. Hashimoto provided medical guidance, data collection, diagnosis, and other insights; and J. Hashimoto and H. Iyatomi were responsible for the study design, overall management, and writing of the paper.

### v5

@@ 5 Results
Figure 5 shows an example of the augmented image obtained with the proposed R-sGAIA for the leftmost original image. Figure 6 compares cancer detection on the EfficientDet-D7 network (baseline model 2) with and without the proposed R-sGAIA and HBBT. The input images are shown in (a), the detection results with RandAugment in (b), and those with the proposed R-sGAIA + HBBT in (c). The proposed system (c) accurately detected the gastric cancer region without being affected by mucosal irregularities or small areas, as demonstrated in the top and middle rows. However, as shown in the bottom row, the result nearly covered the ROI but was slightly misaligned, which led to a false-positive evaluation.
The detection performance of gastric cancer regions from gastric X-ray images is summarized in Table 2. The object detection model detects regions where the probability (i.e., the confidence score) of belonging to the class of interest exceeds a predefined detection threshold of the bounding box $\xi$ as locations containing objects of that class. HBBT assigns new negative labels to the over-detected regions that exceed the threshold $\xi$ and retrains the model. Therefore, to draw an ROC curve for evaluation, the model must be trained every time $\xi$ is changed, which is expensive. In this experiment, therefore, the detection threshold $\xi$ was determined such that the SE (recall) of detecting gastric cancer was about 90%, which was higher than the doctor’s 85.5%. At this detection threshold, HBBT improved precision by 2.3 points in Faster R-CNN (baseline model 1) and 3.3 points in EfficientDet-D7 (baseline model 2). Our system, using the proposed techniques, R-sGAIA and HBBT, achieved a recall of 90.2%, a precision of 42.5%, and an F1 score of 0.578 (Table 2).
The combination of the proposed R-sGAIA and HBBT showed a significant improvement over RandAugment, one of the most advanced data augmentations, in the detection of gastric cancer using the general object detection model. Specifically, the improvement was 10.3 points and 6.1 points in precision, and 10.9 points and 5.9 points in F1 scores for the Faster R-CNN (model 1) and the EfficientDet-D7 (model 2), respectively.
The processing time required for one X-ray image was approximately 0.51 seconds, which makes the proposed system a practical and efficient screening tool for gastric X-ray images.
Figure 6: Example of cancer detection on EfficientDet-D7 networks (as shown in red box with “malignant” label): (a) ground-truth region, (b) results obtained with RandAugment (baseline model 2), and (c) results with R-sGAIA + HBBT (proposed). The proposed R-sGAIA + HBBT led to the accurate detection for the top and middle examples, whereas for the bottom result, although it showed some improvement, the detection was inaccurate.
@@
Augmentation HBBT Precision Recall F1 score
RandAug sGAIA R-sGAIA (%) (%)
@@ 6 Discussion
The cancer detection performance of the proposed system in this study was significantly better than that achieved using baseline methods, including RandAugment (a state-of-the-art general data augmentation strategy), existing methods such as sGAIA (Okamoto et al., 2019), and those proposed by Laddha et al. (2019). Additionally, the system’s cancer detection capability exceeded that of a physician’s reading ($\sim$85.5%). In more than two out of five cases, the detection results presented by the system were cancerous. For a physician using the system, this means that by simply checking the areas surrounded by the few bounding boxes presented in a gastric X-ray image, a higher diagnostic yield can be achieved than with a conventional reading. In other words, the system is practically effective for on-site medical examinations and can significantly reduce the burden on physicians.
@@ 6.1 R-sGAIA
The proposed R-sGAIA can probabilistically highlight the gastric folds on a pixel-by-pixel basis based on medical findings, as shown in Figure 5. In this study, we found that R-sGAIA achieved an F1 score improvement of 6.0 points and 2.7 points over RandAugment on baseline models 1 and 2, respectively. This highlights the importance of probabilistic augmentation informed by medical knowledge. In baseline model 2, R-sGAIA further improved the F1 score by 1.2 points over the previous sGAIA in a direct comparison. This improvement may be attributed to R-sGAIA’s use of validation data to determine $p(e)$ as a continuous function, which enables it to generate a wider variety of enhancement images that better highlighted the gastric folds compared to sGAIA, which used a subjectively fixed value. The function of $p(e)$ with respect to edge strength $e=E(x,y)$ ranged approximately between 0.1 and 0.9, which likely contributed to the performance improvement. By allowing more stochastic variation in regions with very strong and weak edge strengths, R-sGAIA enhanced the diversity of the generated images, thus leading to better detection performance.
@@ 6.2 HBBT
Recent state-of-the-art object detection models (e.g., YOLOv10 (yolov10)) are trained to minimize the difference between the boxes specified in the training data and the currently detected boxes, both in terms of their classes (i.e., contents) and locations. In other words, the model’s error is the sum of the class error and position error for each box, and training is performed to minimize these errors.
The proposed HBBT introduces a new and effective strategy that allows learning from healthy control images, which cannot be used to train conventional object detection models. It retrains over-detected boxes by assigning them a unique “hard sample” class label to suppress future false detections. The HBBT strategy increases the class error for error-prone regions (i.e., hard samples) and retrains the model to reduce false positives. In essence, the HBBT learning algorithm is a strategy to reduce generic errors in object detection models, rather than focusing on the specific type of model. Furthermore, HBBT actively incorporates negative samples, such as images of healthy controls, which are typically excluded from training by conventional object detection models.
Although HBBT is a simple method that recursively learns to eliminate false positives detected in images, including control images, it alone improved precision by 2.3 points and 3.3 points as well as F1 scores by 4.9 points and 2.9 points in baseline models 1 and 2, respectively. HBBT is a general-purpose learning method that can be combined with any object detection algorithm, not just for gastric cancer detection, as demonstrated in this study. Because the final model detection is determined by the bounding box confidence threshold, HBBT not only reduces the number of false positives but also helps minimize false negatives (as seen in the middle case in Figure 6).
6.3 For practical diagnostic support
In this diagnostic support system, the ability to handle the diversity in size and shape of the tumor area to be detected depends on the detectability of the object detection model used. The range of object sizes that can be detected is determined by hyperparameters, such as anchor size and aspect ratio. In the task of cancer detection from gastric X-ray images, the general settings of the object recognition model are sufficient, as the images are taken in a way that ensures the tumor size is large enough for visual diagnosis by a physician. However, for extremely small tumors, it is necessary to adjust the hyperparameters to account for their size.
@@ 6.3 For practical diagnostic support
Detecting gastric cancer regions from X-ray images is inherently challenging, not only for automatic diagnostic systems but also for experts. Despite these challenges, our two technological innovations, R-sGAIA and HBBT, significantly improved the detection performance of gastric cancer, thereby achieving a level of performance and speed that can be effectively used by physicians as a diagnostic aid. Our diagnostic support system offers a concrete solution to the challenges of gastric radiography, in which it is easy to acquire images but difficult to diagnose. In the future, we aim to work closely with experts to further refine and improve the system.
@@ 6.4 Comparison with automated diagnostic support methods for other modalities
Many diagnostic support studies have focused on endoscopy, the most commonly used method for diagnosing gastric cancer. However, as mentioned earlier, most of these studies did not properly separate the training data from the evaluation data, which led to data leakage and artificially inflated diagnostic performance. Recent studies that have addressed this issue by using more sophisticated machine learning methods (Shibata et al., 2020; Ahmad et al., 2023) have reported detection performance for gastric cancer with an image-based F1 score of approximately 0.7, which is considered the current standard for diagnostic accuracy.
Due to its noninvasiveness and ability to image the entire body, X-ray computed tomography (CT) is widely used for detecting not only gastric cancer but also other organs and diseases. Diagnostic studies of gastric cancer using CT typically rely on classical radiomics, including necessary pre-treatment steps. Depending on the cancer stage, diagnostic performance usually achieves an AUC of 0.7 to 0.8 (Meng et al., 2020; Huang et al., 2022). The application of deep learning in this context primarily focuses on predicting patient prognosis and segmenting tumor areas rather than directly diagnosing cancer. Hao et al. (2022) demonstrated that radiomic features and machine learning-extracted features (i.e., low-dimensional representations) are effective at predicting the prognosis of gastric cancer patients, which indicates that these techniques will become more prevalent in the future.
While the diagnostic performance of these automated support methods may appear lower than the 95% SE reported by physicians for gastric cancer, it is essential to note that this comparison is patient-based which makes direct comparisons challenging. As automated diagnostic methods are introduced into clinical practice, their actual performance is likely to approach the current SE of physicians, particularly when multiple outputs are integrated to inform a final decision and provide significant support to physicians. Although our F1 score of 0.578 for the diagnostic performance of gastric X-ray images may seem lower than that of other modalities, it is important to recognize that this is a box-based evaluation. Additionally, because physicians diagnose gastric cancer by reviewing multiple images during a gastric X-ray examination, our method’s ability to highlight the ROI in each image makes it highly valuable. Furthermore, our approach is significant as it represents the first use of gastric X-ray images to aid in the diagnosis of gastric cancer.
@@ 6.5 Limitations of the study
In recent years, it has become apparent that medical images, such as magnetic resonance (MR) images, contain features unique to each facility due to differences in imaging equipment, protocols, geometry, and other factors. These so-called domain differences can significantly impact the results of machine-learning tasks, such as classification, when applied without adjustment. The gastric X-ray images used in this study were only those taken at medical facilities affiliated with Tokai University, and the number of patients (145) and images (4,724) was limited, so the diversity that could be considered was limited. Differences between institutions and the potential lack of diversity in the training data may affect the analysis of gastric X-ray images taken at different institutions, but this was not verified in this study. Finally, this experiment did not allow for a detailed discussion of the detection threshold for hard boundary boxes. Addressing this will be an important focus for future work.
@@ 7 Conclusions
In this paper, we have proposed an unprecedented, accurate, and fast gastric cancer diagnostic support system from gastric X-ray images built on two key innovations: (1) R-sGAIA and (2) HBBT. The proposed system (R-sGAIA + HBBT on the EfficientDet-D7 network) achieves an SE of 90.2%, surpassing that of physicians (85.5%), with a processing time of approximately 0.5 seconds per image. Notably, 42.5% of the detected candidate box regions contained cancer, meaning our system effectively narrows down the areas that physicians need to examine. This system also achieves an F1 score 5.9 points higher than the same network using RandAugment, a state-of-the-art data augmentation method. Diagnosing gastric cancer using gastric X-rays is a challenging task that requires extensive experience. We hope that our proposal will assist in diagnosis and lead to early detection and a reduction in the number of deaths from gastric cancer in the near future.
@@ Acknowledgments
H. Okamoto was responsible for the practical system development of the study and the writing of the paper; Q. H. Cap provided model implementation, code publication, and paper writing; T. Nomura, K. Nabeshima, and J. Hashimoto provided medical guidance, data collection, diagnosis, and other insights; and J. Hashimoto and H. Iyatomi were responsible for the study design, overall management, and writing of the paper.